OpenStudy (anonymous):

What is the volume of a cube with side length twelve feet? seventy-two cubic feet one hundred forty-four cubic feet eight hundred sixty-four cubic feet one thousand seven hundred twenty-eight cubic feet

OpenStudy (anonymous):

Okay the volume of a cube is found using this formula: v=s^3 s^3 is the same as s x s x s where s is the length of one of the sides

OpenStudy (anonymous):

sorry bu what does ^ mean? >.<

OpenStudy (anonymous):

it means to a power, sorry do you know what happens if you square something?

OpenStudy (anonymous):

No,

OpenStudy (anonymous):

err maybe

OpenStudy (anonymous):

Okay when you square something, you multiply the number by itself

OpenStudy (anonymous):

oh so like 311. it would be 311 x 311?

OpenStudy (anonymous):

or 311 x 1

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

k

OpenStudy (anonymous):

nono you had it right, but since it is asking for l*w*h to find volume, we must multiply the side length times itself 2 times instead of just once

OpenStudy (anonymous):

so s^3 would be s = 3?

OpenStudy (anonymous):

no, s^3 is s*s*s where s is the length of one side

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Since 12 is the length of the side, s=12

OpenStudy (anonymous):

12 x 3

OpenStudy (anonymous):

Not exactly, remember I said it is a number multiplied to itself?

OpenStudy (anonymous):

would it be 12x12 or 3x3?

OpenStudy (anonymous):

It is the same thing you did before like 311*311 just after you calculate that, you multiply it again

OpenStudy (anonymous):

im thinking 12?...

OpenStudy (anonymous):

Yes 12 is right since s=12

OpenStudy (anonymous):

12x12 that is

OpenStudy (anonymous):

So after you do 12*12 you multiply that answer *12 again and that is the volume

OpenStudy (anonymous):

so 12x12 is 144

OpenStudy (anonymous):

144 x 12 is1728

OpenStudy (anonymous):

That is one of your answer choices

OpenStudy (anonymous):

164 cubic feet

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

is it the right one? or should i calculate again? 1728

OpenStudy (anonymous):

V=12^3=1728 he is right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

tks

OpenStudy (anonymous):

No problem